Assume that Country A has a population of 500,000 and only produces 1 good: cars. Country A produces 100,000 cars per year. The people in Country A purchase 90,000 cars, but there are not enough cars to fulfill all the demand. They decide to import 50,000 more. The government buys 25,000 cars for its police force, and 10,000 cars are bought by companies to transport employees to other locations to work. They also export 65,000 cars to nearby countries for sale. Discuss the following:
• What is Country A’s GDP?
• What is the composition of GDP by percentage?
• What is the GDP per capita?
• How does this relate to Keynesian economics?
1
Expert's answer
2014-10-28T14:32:07-0400
Population of 500,000, produce 100,000 cars per year, people purchase 90,000 cars, imports of 50,000 more, government buys 25,000 cars for its police force, 10,000 cars are bought by companies, exports of 65,000 cars.
• Country A’s GDP is: GDP = Consumption (C) + Investment (I) + Government purchases (G) + Net export (EX - IM) = 90,000 + 10,000 + 25,000 + (65,000 - 50,000) = 140,000 cars.
• The composition of GDP by percentage is: consumption - 64%, investment - 7%, government purchases - 18%, net exports - 11%.
• The GDP per capita is GDP/Population = 140,000/500,000 = 0.28 car per person.
• Keynesian economics is the view that in the short run, especially during recessions, economic output is strongly influenced by aggregate demand (total spending in the economy). In the Keynesian view, aggregate demand does not necessarily equal the productive capacity of the economy; instead, it is influenced by a host of factors and sometimes behaves erratically, affecting production, employment, and inflation.